OpenStudy (stevetron):

factor the expression. 4x^2-81y^2

OpenStudy (haseeb96):

a^2 - b^2= (a-b)(a+b) 4x^2 -81y^2=(2x -9y)(2x +9y) is your answer

OpenStudy (stevetron):

oh OOOHHHHHHH, ok thank you so much!

OpenStudy (stevetron):

oh OOOHHHHHHH, ok thank you so much!